OpenStudy (imnotarobot5):

Find f '(x) for f(x)= sqrt(sin (4x))?

ganeshie8 (ganeshie8):

Can you find the derivative of sqrt(x) ?

OpenStudy (agent0smith):

Re-writing the square root as a 1/2 exponent may help. And you need to be familiar with the chain rule.

OpenStudy (trojanpoem):

f'(x) = 4Cos4x/2sqrt(sin4x) using the formula f'(x) = dx/2sqrt(x) or take the long route. f'(x) = 0.5 (sin4x)^(-0.5) * 4(cos4x) = 4cos4x/2sqrt(sin4x)